1)     What is the mass of water if it contains 5 moles of hydrogen atoms?                              

2)      Aspirin is a compound with the molecular formula C9H8O4. What is its percent composition?

3)     A 24.81-g sample of a gaseous compound containing only carbon, oxygen, and chlorine is determined to contain 3.01 g C, 4.00 g O, and 17.81 g Cl. What is this compound’s percent composition?


1
Expert's answer
2022-02-17T13:17:02-0500

1) Molar mass of water = 18.0g

Mass"= \frac{5}{2}×18=45g"


2)C= 12×9= 108g

H= 1×8=8g

O=16×4=64g

Total= 180g


C= "\frac{108}{180}×100\%=60\%"


H"= \frac{8}{180}×100\%=4.44\%"


O="\frac{64}{180}×100\%=35.56\%"


3) "24.81\ne3.01+4.00+17.81"

Taking 3.01 as erroneous and using 3.00 instead

"\frac{3.00}{24.81}×100\%=12.09\%"


"\frac{4.00}{24.81}×100\%=16.12\%"


"\frac{17.81}{24.81}×100\%=71.79\%"